The diameter AB is equal to the sum of the six smaller diameters. What is the differences in the number of the units between the circumference of the large circle and the combined circumferences of the six smaller circles? Explain in detal.
Answer by ankor@dixie-net.com(22740)

The diameter AB is equal to the sum of the six smaller diameters. What is the differences in the number of the units between the circumference of the large circle and the combined circumferences of the six smaller circles? Explain in detail.
:
Let 6d = the diameter of large circle
:
Then d = the diameter of the small circles
:
Large circle circum = pi*6d
Small circle circum = pi*d
:
Total circum of 6 small circles = 6(pi*d) = 6pi*d. same as pi*6d
:
Did this help?